OpenStudy (anonymous):

Solve the following inequality and write your answer using interval notation. (x+4/10)≤(2/11-3x)

OpenStudy (anonymous):

(x+4)/10 - 2/(11-3x) <= 0 multiplu yhtu by 10(11-3x) (x+4)(11-3x) - 20 <= 0 11 - 3x^2 + 44 -12x - 20 >= 0 3x^2 + x -24 <= 0 (3x-8)(x+3) <= 0 critical points are 8/3 and -3 graph U shaped solution is x <= -3 , x >= 8/3

OpenStudy (anonymous):

oh you need interval notation x = (-infinity,-3] and [8/3, +infinity)

OpenStudy (anonymous):

theres a typo on 4th line - 11x not 11

OpenStudy (radar):

Jimmy, I am having following your 4th line 11-3x^2+44-12x-20 how does this become 3x^2 +x-24?

OpenStudy (anonymous):

11x and also another typo <= not >= - sorry

OpenStudy (radar):

I see

OpenStudy (anonymous):

i did it ok on paper but made typos when copying i think its correct bu ill check it out on wolfram alpha

OpenStudy (radar):

I got down to -3x^2-x+44<=0 3x^2 +x _44>=0 and got confused!

OpenStudy (radar):

What did Wolfram alpha say?

OpenStudy (anonymous):

what happened to the rest of the post??????

OpenStudy (radar):

I see that Wolfram Alpha got 11/3 on the positive side.

OpenStudy (anonymous):

the correct solution according to wolfram is x <= -3 as i got or (-infinity,-3] and also [8/3, 11/3) i somehow missed another critical point in my solution.

OpenStudy (radar):

(x+4)(3x-11)>=0 As you can see I messed up somewhere.

OpenStudy (radar):

There is a difference between -4 and -8/3 !!lol

OpenStudy (radar):

Hey between us, we got the right points. lol

OpenStudy (radar):

mandaboo, please be tolerant lol

OpenStudy (jamesj):

@jimmy: the problem is in your early step when you multiplied through by 11-3x We have to consider the two cases where this is positive and where it is negative.

OpenStudy (anonymous):

aahh - of course ! - thanx james